# Smoother-type a posteriori error estimates for finite element methods

### v2

@@ 3.2 A Posteriori Error Estimates in H(curl)
Smoother-type a posteriori error estimation could be generalized to vector-valued finite elements. We consider the problem: find $u\in H({\rm curl},\Omega)$ and $u_{h}\in V_{h}\subset H({\rm curl},\Omega)$
@@
where $V_{h}(\rm curl)$ is a Nédélec edge element space, and $H({\rm curl},\Omega)=\{v\in L^{2}(\Omega):{\rm curl}v\in L^{2}(\Omega)\}$ is equipped with the norm $\|\bullet\|_{H(\rm curl,\Omega)}=\big(\|\bullet\|_{L^{2}(\Omega)}^{2}+\|{\rm curl}\bullet\|_{L^{2}(\Omega)}^{2}\big)^{\frac{1}{2}}$. Here ${\rm curl}v=(\frac{\partial v_{3}}{\partial x_{2}}-\frac{\partial v_{2}}{\partial x_{3}},\frac{\partial v_{1}}{\partial x_{3}}-\frac{\partial v_{3}}{\partial x_{1}},\frac{\partial v_{2}}{\partial x_{1}}-\frac{\partial v_{2}}{\partial x_{2}})$ in $\mathbb{R}^{3}$ and ${\rm curl}v=\frac{\partial v_{2}}{\partial x_{1}}-\frac{\partial v_{1}}{\partial x_{2}}$ in $\mathbb{R}^{2}$.
@@
is a uniform preconditioner for $A_{h/2}({\rm curl})$. As a result, the block smoother $S_{a}=S_{a}({\rm curl})$ in Theorem 2.1 leads to uniform a posteriori error estimate
$\|u-u_{h}\|_{H({\rm curl},\Omega)}\eqsim\langle r,S_{a}({\rm curl})r\rangle^{\frac{1}{2}}=\Big(\sum_{k=1}^{N_{h/2}}\|\eta_{k}\|_{H({\rm curl},\Omega_{h/2,k})}\Big)^{\frac{1}{2}}.$ (3.9)
@@
The H(curl) problem (3.8) is discretized by the lowest order edge finite element on the L-shaped domain $\Omega=(-1,1)^{2}\backslash([0,1)\times[-1,0))$ as well as on the 3D cavity region $\Omega=(0,1)^{3}\backslash[0.25,0.75]^{3}$, see Figure 3.7. We compare the smoother-type a posteriori error estimates (3.9) and (3.10) with the residual-type error estimator (cf. [16, 60, 52])
@@
Here $n\wedge{\rm curl}u_{h}|_{F}=t\cdot{\rm curl}u_{h}$ in $\mathbb{R}^{2}$ with $t$ being the unit tangent vector along $F$ and $n\wedge{\rm curl}u_{h}|_{F}=n\times{\rm curl}u_{h}$ in $\mathbb{R}^{3}$ with $n$ being the unit normal to $F$. On the L-shaped domain $\Omega$, the exact solution is set to be $u=\nabla(r^{2/3}\sin(2\theta/3))$, where $r^{2/3}\sin(2\theta/3)$ is the singular solution used in Section 3.1. On the 3D test region, we set $f=(\sqrt{x^{2}-2y^{2}+3z^{2}+3}\exp(-xyz),0,0)$ in (3.8) and use $\|u_{h}-u_{h/2}\|_{H(\rm curl,\Omega)}$ as the reference FEM error as the exact solution $u$ is unknown. It is shown in Figure 3.5 that all estimators lead to optimally convergent FEM solutions, while smoother-type a posteriori error estimates are much closer to the exact error than the residual error estimators.
@@ 4 Block Diagonal Estimators for Symmetric Systems
$\langle r_{\sigma},\tau\rangle$ $:=g(\tau)-a(\sigma_{h},\tau)-b(\tau,u_{h}),\quad\tau\in\Sigma_{h/2}^{*},$
$\langle r_{u},v\rangle$ $:=f(v)-b(\sigma_{h},v),\quad v\in U_{h/2}^{*}.$
@@
$\|(\sigma_{h/2}-\sigma_{h},u_{h/2}-u_{h})\|_{\Sigma\times U}\leq\beta_{h}^{-1}\|(r_{\sigma},r_{u})\|_{\Sigma^{*}\times U^{*}}$
$=\beta_{h}^{-1}\sqrt{\langle r_{\sigma},A_{\Sigma_{h/2}}^{-1}r_{\sigma}\rangle+\langle r_{u},A^{-1}_{U_{h/2}}r_{u}\rangle}$
$\leq\beta_{h}^{-1}\sqrt{c_{3}\langle r_{\sigma},B_{\Sigma_{h/2}}r_{\sigma}\rangle+c_{5}\langle r_{u},B_{U_{h/2}}r_{u}\rangle},$
@@
$\|(\sigma_{h/2}-\sigma_{h},u_{h/2}-u_{h})\|_{\Sigma\times U}\leq\frac{\sqrt{\max(c_{3},c_{5})}}{\beta_{h}}\sqrt{\langle r_{\sigma},S_{\Sigma_{h/2}}r_{\sigma}\rangle+\langle r_{u},S_{U_{h/2}}r_{u}\rangle}.$ (4.4)
@@
$\|(\sigma-\sigma_{h},u-u_{h})\|_{\Sigma\times U}\leq\frac{\sqrt{\max(c_{3},c_{5})}}{\beta_{h}(1-\gamma_{h/2})}\sqrt{\langle r_{\sigma},S_{\Sigma_{h/2}}r_{\sigma}\rangle+\langle r_{u},S_{U_{h/2}}r_{u}\rangle}.$
@@ 4.1 Poisson in Mixed Form
Let $\Sigma_{h}=V_{h}({\rm div})$ be the lowest order Raviart-Thomas finite element space equipped with the norm $\|\bullet\|_{H(\rm div,\Omega)}=\big(\|\bullet\|_{L^{2}(\Omega)}^{2}+\|{\rm div}\bullet\|_{L^{2}(\Omega)}^{2}\big)^{\frac{1}{2}}$ and inner product $\langle A_{h}({\rm div})\bullet,\bullet\rangle$. Let $U_{h}$ be the space of piecewise constants on $\mathcal{T}_{h}$ equipped with the $L^{2}(\Omega)$ norm. The mixed FEM for (2.8) seeks $(\sigma_{h},u_{h})\in V_{h}({\rm div})\times U_{h}$ such that
$(\sigma_{h},\tau_{h})+({\rm div}\tau_{h},u_{h})$ $=0,\quad\tau_{h}\in V_{h}({\rm div}),$ (4.5)
@@
Let $S_{a}({\rm div}):V_{h/2}({\rm div})^{*}\rightarrow V_{h/2}({\rm div})$ be the additive block smoother on $\mathcal{T}_{h/2}$, i.e., $S_{a}({\rm div})r=\sum_{k=1}^{N_{h/2}}\eta_{k}$ with each $\eta_{k}\in V_{k}({\rm div})$ solving
$({\rm div}\eta_{k},{\rm div}\tau)_{\Omega_{h/2,k}}+(\eta_{k},\tau)_{\Omega_{h/2,k}}$
$=\langle r_{\sigma},v\rangle=-(\sigma_{h},\tau)_{\Omega_{h/2,k}}-({\rm div}\tau,u_{h})_{\Omega_{h/2,k}},\quad\tau\in V_{k}({\rm div}).$
@@
$\|\sigma-\sigma_{h}\|^{2}_{H({\rm div},\Omega)}+\|u-u_{h}\|^{2}_{L^{2}(\Omega)}\eqsim\langle r_{\sigma},S_{a}({\rm div})r_{\sigma}\rangle+\langle r_{u},A_{U_{h/2}}^{-1}r_{u}\rangle.$ (4.6)
@@
For the mixed FEM (4.5), we compare the smoother-type a posteriori error estimate (4.6) to the residual error estimator (cf. [22, 25, 48])
@@
$\quad+\sum_{E\text{ is an edge in }\mathcal{T}_{h}}h_{E}\|\llbracket\sigma_{h}\cdot t\rrbracket\|^{2}_{L^{2}(E)}+h_{E}\|\llbracket u_{h}\rrbracket\|^{2}_{L^{2}(E)}\Big)^{\frac{1}{2}}.$
Figure 4.1 demonstrates that the smoother-type estimator (4.6) is more accurate than $\mathcal{E}_{\rm res,mix}$ and yields an optimally convergent AFEM.
@@ 4.2 Biot’s Consolidation Model
$\|q_{T}\|_{Q_{T}}$ $=\langle A_{Q_{T}}v,v\rangle^{\frac{1}{2}}:=\sqrt{\lambda^{-1}(q_{T},q_{T})+(\bar{q}_{T},\bar{q}_{T})},$
@@ 5.1 Convection-diffusion equation
$-\nabla\cdot(\alpha\nabla u+\beta u)$ $=f\quad\text{ in }\partial\Omega,$ (5.1)
@@
The experimental results in Figure 5.2 demonstrate that the smoother-type estimators remain effective for the highly non-symmetric problem (5.1). As shown in Figure 5.1, the corresponding AFEM is able to capture both the interface and boundary singularity. It is also observed from Figure 5.2 that the effectiveness ratios $|S^{a}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}/|u-u_{h}|_{H^{1}(\Omega)}$ and $|S^{m}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}/|u-u_{h}|_{H^{1}(\Omega)}$ remain mild as the polynomial degree $p$ grows.
@@ 5.2 Helmholtz Equation
Here $k>0$ denotes the wavenumber, $\texttt{i}=\sqrt{-1}$ is the imaginary unit, and (5.2b) is an approximation to the Sommerfeld radiation condition at infinity. For large wavenumber $k\gg 1$, the solution of (5.2) is highly oscillatory, leading to challenges in numerical approximation due to increased computational cost and pollution effects. Let $H_{0}^{1}(\Omega,\mathbb{C})$ be the complex-valued Sobolev space and $V_{h}\subset H_{0}^{1}(\Omega,\mathbb{C})$ be the continuous and piecewise $\mathcal{P}_{p}$ finite element subspace. The FEM for (5.2) seeks $u_{h}\in V_{h}$ such that
@@
In the numerical experiment, we set $\Omega_{1}=(-0.5,0.8)\times(-0.5,0.5)$ and $\Omega_{0}$ as a triangle with vertices $(0,0)$, $(0.5,\pm 0.5\tan(\theta_{0}))$ with $\theta_{0}=\pi/6$. Let $(r,\theta)$ be the polar coordinate near the origin and $\alpha=\pi/(2\pi-2\theta_{0})$. The exact solution of (5.2) is $u=\phi(r)J_{\alpha}(kr)\sin\left(\alpha(\theta-\theta_{0})\right)$, where $J_{\alpha}$ is the Bessel function of the first kind and
@@
is a cutoff function. The polynomial degree $p$ ranges from 1 to 7 and the initial mesh size $h_{0}$ is $h_{0}\approx 1/60$ for $p=1$ and $h_{0}\approx 1/30$ for $p\geq 2$. The experimental results in Figure 5.5 demonstrate that the smoother-type estimators remain effective for the Helmholtz equation with large wave number $k=40\pi$. As shown in Figure 5.4, AFEMs based on smoother-type a posteriori error estimates correctly capture the oscillation and corner singularity of (5.2). It is also observed from Figure 5.5 that the effectiveness ratios $|S^{a}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}/|u-u_{h}|_{1,k}$ and $|S^{m}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}/|u-u_{h}|_{1,k}$ are similar and uniform as the polynomial degree $p$ grows. Considering the computational cost, $|S^{a}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}$ is more efficient than its Gauss–Seidel counterpart.
Figure 5.4: Adaptive meshes (approximately $10^{5}$ triangles) by $\mathcal{P}_{1}$-AFEM (left) and $\mathcal{P}_{4}$-AFEM (right) driven by the pointwise Jacobi smoother error estimator $|S^{a}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}$.

### v3

@@ 3.2 A Posteriori Error Estimates in H(curl)
Smoother-type a posteriori error estimation could be generalized to vector-valued finite elements. In $\mathbb{R}^{d}$ with $d=2$ or 3, we consider the problem: find $u\in H({\rm curl},\Omega)$ and $u_{h}\in V_{h}\subset H({\rm curl},\Omega)$
@@
where $V_{h}(\rm curl)$ is a Nédélec edge element space, and $H({\rm curl},\Omega)=\{v\in[L^{2}(\Omega)]^{d}:{\rm curl}v\in[L^{2}(\Omega)]^{d(d-1)/2}\}$ is equipped with the norm $\|\bullet\|_{H(\rm curl,\Omega)}=\big(\|\bullet\|_{L^{2}(\Omega)}^{2}+\|{\rm curl}\bullet\|_{L^{2}(\Omega)}^{2}\big)^{\frac{1}{2}}$. Here ${\rm curl}v=(\frac{\partial v_{3}}{\partial x_{2}}-\frac{\partial v_{2}}{\partial x_{3}},\frac{\partial v_{1}}{\partial x_{3}}-\frac{\partial v_{3}}{\partial x_{1}},\frac{\partial v_{2}}{\partial x_{1}}-\frac{\partial v_{2}}{\partial x_{2}})$ in $\mathbb{R}^{3}$ and ${\rm curl}v=\frac{\partial v_{2}}{\partial x_{1}}-\frac{\partial v_{1}}{\partial x_{2}}$ in $\mathbb{R}^{2}$.
@@
is a uniform preconditioner for $A_{h/2}({\rm curl})$. As a result, the block smoother $S_{a}=S_{a}({\rm curl})$ in Theorem 2.1 leads to a uniform a posteriori error estimate
$\|u-u_{h}\|_{H({\rm curl},\Omega)}\eqsim\langle r,S_{a}({\rm curl})r\rangle^{\frac{1}{2}}=\Big(\sum_{k=1}^{N_{h/2}}\|\eta_{k}\|^{2}_{H({\rm curl},\Omega_{h/2,k})}\Big)^{\frac{1}{2}}.$ (3.9)
@@
The $H(\rm{curl})$ problem (3.8) is discretized by the lowest order edge finite element on the L-shaped domain $\Omega=(-1,1)^{2}\backslash([0,1)\times[-1,0))$ as well as on the 3D cavity region $\Omega=(0,1)^{3}\backslash[0.25,0.75]^{3}$, see Figure 3.7. We compare the smoother-type a posteriori error estimates (3.9) and (3.10) with the residual-type error estimator (cf. [16, 64, 54])
@@
Here $n\wedge{\rm curl}u_{h}|_{F}=t\cdot{\rm curl}u_{h}$ in $\mathbb{R}^{2}$ with $t$ being a unit tangent vector along $F$ and $n\wedge{\rm curl}u_{h}|_{F}=n\times{\rm curl}u_{h}$ in $\mathbb{R}^{3}$ with $n$ being the unit normal to $F$. On the L-shaped domain $\Omega$, the exact solution is set to be $u=\nabla(r^{2/3}\sin(2\theta/3))$, where $r^{2/3}\sin(2\theta/3)$ is the singular solution used in Section 3.1. On the 3D test region, we set $f=(\sqrt{x^{2}-2y^{2}+3z^{2}+3}\exp(-xyz),0,0)$ in (3.8) and use $\|u_{h}-u_{h/2}\|_{H(\rm curl,\Omega)}$ as the reference FEM error as the exact solution $u$ is unknown. It is shown in Figure 3.5 that all estimators lead to optimally convergent FEM solutions, while smoother-type a posteriori error estimates are much closer to the exact error than the residual error estimators.
@@ 4 Block Diagonal Estimators for Symmetric Systems
$\langle r_{\sigma},\tau\rangle$ $:=g(\tau)-a(\sigma_{h},\tau)-b(\tau,u_{h}),\quad\tau\in\Sigma_{h/2},$
$\langle r_{u},v\rangle$ $:=f(v)-b(\sigma_{h},v),\quad v\in U_{h/2}.$
@@
$\|(\sigma_{h/2}-\sigma_{h},u_{h/2}-u_{h})\|_{\Sigma\times U}\leq{\beta_{h/2}^{-1}}\|(r_{\sigma},r_{u})\|_{\Sigma^{*}\times U^{*}}$
$=\beta_{h/2}^{-1}\sqrt{\langle r_{\sigma},A_{\Sigma_{h/2}}^{-1}r_{\sigma}\rangle+\langle r_{u},A^{-1}_{U_{h/2}}r_{u}\rangle}$
$\leq{\beta_{h/2}^{-1}}\sqrt{c_{3}\langle r_{\sigma},B_{\Sigma_{h/2}}r_{\sigma}\rangle+c_{5}\langle r_{u},B_{U_{h/2}}r_{u}\rangle},$
@@
$\|(\sigma_{h/2}-\sigma_{h},u_{h/2}-u_{h})\|_{\Sigma\times U}\leq\frac{\sqrt{\max(c_{3},c_{5})}}{\beta_{h/2}}\sqrt{\langle r_{\sigma},S_{\Sigma_{h/2}}r_{\sigma}\rangle+\langle r_{u},S_{U_{h/2}}r_{u}\rangle}.$ (4.4)
@@
$\|(\sigma-\sigma_{h},u-u_{h})\|_{\Sigma\times U}\leq\frac{\sqrt{\max(c_{3},c_{5})}}{\beta_{h/2}(1-\gamma_{h/2})}\sqrt{\langle r_{\sigma},S_{\Sigma_{h/2}}r_{\sigma}\rangle+\langle r_{u},S_{U_{h/2}}r_{u}\rangle}.$
@@ 4.1 Poisson in Mixed Form
Let $\Sigma_{h}=V_{h}({\rm div})$ be the lowest order Raviart-Thomas finite element space equipped with the norm $\|\bullet\|_{H(\rm div,\Omega)}=\big(\|\bullet\|_{L^{2}(\Omega)}^{2}+\|{\rm div}\bullet\|_{L^{2}(\Omega)}^{2}\big)^{\frac{1}{2}}$ and inner product $\langle A_{h}({\rm div})\bullet,\bullet\rangle$. Let $U_{h}$ be the space of piecewise constants on $\mathcal{T}_{h}$ equipped with the $L^{2}(\Omega)$ norm. For the boundary value problem $\Delta u=f$ in $\Omega$ with $u|_{\partial\Omega}=g$, the mixed FEM seeks $(\sigma_{h},u_{h})\in V_{h}({\rm div})\times U_{h}$ such that
$(\sigma_{h},\tau_{h})+({\rm div}\tau_{h},u_{h})$ $={(g,\tau_{h}\cdot n)_{\partial\Omega}},\quad\tau_{h}\in V_{h}({\rm div}),$ (4.5)
@@
Let $S_{a}({\rm div}):V_{h/2}({\rm div})^{*}\rightarrow V_{h/2}({\rm div})$ be the additive block smoother on $\mathcal{T}_{h/2}$, i.e., $S_{a}({\rm div})r_{\sigma}=\sum_{k=1}^{N_{h/2}}\eta_{k}$ with each $\eta_{k}\in V_{k}({\rm div})$ solving
$({\rm div}\eta_{k},{\rm div}\tau)_{\Omega_{h/2,k}}+(\eta_{k},\tau)_{\Omega_{h/2,k}}=\langle r_{\sigma},\tau\rangle$
$={(g,\tau\cdot n)_{\partial\Omega\cap\partial\Omega_{h/2,k}}}-(\sigma_{h},\tau)_{\Omega_{h/2,k}}-({\rm div}\tau,u_{h})_{\Omega_{h/2,k}},\quad\tau\in V_{k}({\rm div}).$
@@
$\|\sigma-\sigma_{h}\|^{2}_{H({\rm div},\Omega)}+\|u-u_{h}\|^{2}_{L^{2}(\Omega)}$ $\eqsim\langle r_{\sigma},S_{a}({\rm div})r_{\sigma}\rangle+\langle r_{u},A_{U_{h/2}}^{-1}r_{u}\rangle$ (4.6)
$=\sum_{k=1}^{N_{h/2}}\|\eta_{k}\|_{H({\rm div},\Omega_{h/2,k})}^{2}+\langle r_{u},A_{U_{h/2}}^{-1}r_{u}\rangle.$
@@
For the mixed FEM (4.5), we compare the smoother-type a posteriori error estimate (4.6) to the residual error estimator (cf. [22, 25, 50, 49])
@@
$+\sum_{\text{edge }E\subset\mathring{\Omega}}h_{E}\|\llbracket\sigma_{h}\cdot t\rrbracket\|^{2}_{L^{2}(E)}+h_{E}\|\llbracket u_{h}\rrbracket\|^{2}_{L^{2}(E)}{+\sum_{\text{edge }E\subset\partial\Omega}h_{E}\|\sigma_{h}\cdot t-\partial g/\partial t\|^{2}_{L^{2}(E)}+h_{E}\|u_{h}-g\|^{2}_{L^{2}(E)}\Big)^{\frac{1}{2}}}.$
In the experiment, we use the same exact singular solution $u$ in Section 3.1.3. Figure 4.1 demonstrates that the smoother-type estimator (4.6) is more accurate than $\mathcal{E}_{\rm res,mix}$ and yields an optimally convergent AFEM.
@@ 4.2 Biot’s Consolidation Model
$\|q_{T}\|_{Q_{T}}$ $=\langle A_{Q_{T}}q_{T},q_{T}\rangle^{\frac{1}{2}}:=\sqrt{\lambda^{-1}(q_{T},q_{T})+(\bar{q}_{T},\bar{q}_{T})},$
@@ 5.1 Convection-diffusion equation
$-\nabla\cdot(\alpha\nabla u+\beta u)$ $=f\quad\text{ in }\Omega,$ (5.1)
@@
The experimental results in Figure 5.2 demonstrate that the smoother-type estimators remain effective for the highly non-symmetric problem (5.1). As shown in Figure 5.1, the corresponding AFEM is able to capture both the interface and boundary singularity. It is also observed from Figure 5.2 that the effectivity ratios $|S^{a}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}/|u-u_{h}|_{H^{1}(\Omega)}$ and $|S^{m}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}/|u-u_{h}|_{H^{1}(\Omega)}$ remain mild as the polynomial degree $p$ grows.
@@ 5.2 Helmholtz Equation
Here $k>0$ denotes the wavenumber, $\texttt{i}=\sqrt{-1}$ is the imaginary unit, and (5.2b) is an approximation to the Sommerfeld radiation condition at infinity. For large wavenumber $k\gg 1$, the solution of (5.2) is highly oscillatory, leading to challenges in numerical approximation due to increased computational cost and pollution effects. Let $H_{\partial\Omega_{0}}^{1}(\Omega,\mathbb{C})=\{v\in H^{1}(\Omega,\mathbb{C}):v|_{\partial\Omega_{0}}=0\}$ be the complex-valued Sobolev space and $V_{h}\subset H_{{\partial\Omega_{0}}}^{1}(\Omega,\mathbb{C})$ be the continuous and piecewise $\mathcal{P}_{p}$ finite element subspace. The FEM for (5.2) seeks $u_{h}\in V_{h}$ such that
@@
In the numerical experiment, we set $\Omega_{1}=(-0.5,0.8)\times(-0.5,0.5)$ and $\Omega_{0}$ as a triangle with vertices $(0,0)$, $(0.5,\pm 0.5\tan(\theta_{0}))$ with $\theta_{0}=\pi/6$. Let $(r,\theta)$ be the polar coordinates near the origin and $\alpha=\pi/(2\pi-2\theta_{0})$. The exact solution of (5.2) is $u=\phi(r)J_{\alpha}(kr)\sin\left(\alpha(\theta-\theta_{0})\right)$, where $J_{\alpha}$ is the Bessel function of the first kind and
@@
is a cutoff function. The polynomial degree $p$ ranges from 1 to 7 and the initial mesh size $h_{0}$ is $h_{0}\approx 1/60$ for $p=1$ and $h_{0}\approx 1/30$ for $p\geq 2$. The experimental results in Figure 5.5 demonstrate that the smoother-type estimators remain effective for the Helmholtz equation with large wave number $k=40\pi$. As shown in Figure 5.4, AFEMs based on smoother-type a posteriori error estimates correctly capture the oscillation and corner singularity of (5.2). It is also observed from Figure 5.5 that the effectivity ratios $\|S^{a}_{\mathcal{P}_{p}}r\|_{1,k}/\|u-u_{h}\|_{1,k}$ and $\|S^{m}_{\mathcal{P}_{p}}r\|_{1,k}/\|u-u_{h}\|_{1,k}$ are similar and uniform as the polynomial degree $p$ grows. Considering the computational cost, $\|S^{a}_{\mathcal{P}_{p}}r\|_{1,k}$ is more efficient than its Gauss–Seidel counterpart.
Figure 5.4: Adaptive meshes (approximately $10^{5}$ triangles) by $\mathcal{P}_{1}$-AFEM (left) and $\mathcal{P}_{4}$-AFEM (right) driven by the pointwise Jacobi smoother error estimator $\|S^{a}_{\mathcal{P}_{p}}r\|_{1,k}$.